OpenStudy (steven513):

The table on the right shows the measured dimensions of a rectangular prism and the minimum and maximum possible dimensions based on the greatest possible error. What is the greatest possible percent error in finding the volume of the​ prism? (Round to the nearest percent as needed.) Length|Width|Height Measured: 10 7 5 Minimum: 9.5 6.5 4.5 Maximum: 10.5 7.5 5.5 I said the greatest possible percent error in finding the volume of the prism is 3%.

OpenStudy (jiteshmeghwal9):

Volume of rectangular prism = W*L*H Where, W=width L=Length H=Height

OpenStudy (jiteshmeghwal9):

maximum Possible percent error =\[\left( \frac{\Delta L}{L} +\frac{ \Delta W}{W}+\frac{\Delta H}{H} \right)\times 100\]

OpenStudy (jiteshmeghwal9):

maximum value of \(\Delta L\)=10.5-10=0.5 maximum value of \(\Delta W\)=7.5-7=0.5 maximum value of \(\Delta H\)=5.5-5=0.5

OpenStudy (jiteshmeghwal9):

maximum possible percent error=\[\left( \frac{0.5}{10}+\frac{0.5}{7}+\frac{0.5}{5} \right)\times 100\]

OpenStudy (jiteshmeghwal9):

\(\approx 22.14%\)

OpenStudy (steven513):

@jiteshmeghwal9 I get 0.2375, meaning 23.75. what I did was the maximum volume-measure volume divided by the measure volume. Giving 0.2375

OpenStudy (jiteshmeghwal9):

\[\left( \frac{\Delta L}{L} +\frac{ \Delta W}{W}+\frac{\Delta H}{H} \right)\times 100\]u must apply this formula to get the percentage error